what is a one digit number and a two digit number whose product is a number that ends in two zeros.
Answer by Edwin McCravy(20086)

There are 12 possible answers:

 1.  2 × 50 = 100
 2.  4 × 25 = 100
 3.  4 × 50 = 200
 4.  4 × 75 = 300
 5.  5 × 20 = 100
 6.  5 × 40 = 200
 7.  5 × 60 = 300
 8.  5 × 80 = 400
 9.  6 × 50 = 300
10.  8 × 25 = 200
11.  8 × 50 = 400
12.  8 × 75 = 600

Edwin